OpenStudy (keana):

add & simplify x/3x+2x-3/3x

OpenStudy (anonymous):

if it is \[\frac{x}{3x}+\frac{2x-3}{3x}\] since the denominators are the same, you can add the numerators \[=\frac{x+2x-3}{3x}\] thne you can combine like terms, \[\frac{3x-3}{3x}\] factor \[\frac{3(x-1)}{3x}\] and finally cancel the 3

OpenStudy (keana):

@satellite73 thanks so much !